OpenStudy (anonymous):

b. Solve this radical equation: sqrt{3x+2}+sqrt{x+7}=1 . (You need to tell Patricia that simply squaring both sides will not do the trick and explain why. Discuss the method used to solve such equations. Show her how once you set up the equation correctly, it’s simply a quadratic equation that can be solved by straightforward methods.)

OpenStudy (anonymous):

\[\sqrt{3x+2}+\sqrt{x+7}=1 \]